Assume that the unit Z measures the level of interest that the average CHEM150 student shows in Dr. Zimmermann's lectures.

Convert 0.0104 kZ into mZ!

Enter the numerical result (in mZ) without the unit.
Answer #1

1 kZ = 10^6 mZ

0.0104 kZ = 10^6*0.0104

= 1.04*10^4 mZ

If you have any query please comment

If you satisfied with the solution please like it thankxx